The length of a rectangle is 3 times the width. The perimeter is 44 cm. What are the dimensions of the rectangle?
Answer by JulietG(1812)

P = 2L +2W
44 = 2L + 2W
L = 3W
Substitute
44 = 2(3W) + 2W
44 = 6W + 2W
44 = 8W
Divide each side by 8
44/8 = W
Reduce
5 1/2 = W
.
If you know the width is 5.5, and the length is 3*5.5, then the length is 16.5.
Prove it by putting in the values:
2(16.5) + 2(5.5) = 44
33 + 11 does indeed equal 44